Put in Standard Form
x^2-10x+y^2+16=0
Cannot figure it out, Thanks

Answer by lwsshak3(11628)

Put in Standard Form
x^2-10x+y^2+16=0
complete the square
(x^2-10x+25)+y^2=-16+25
(x-5)^2+y^2=9
This is a circle with center at (5,0) and radius=3